The weekly demand for Kelewele among the 2018 batch of MBA students at UPSA is Qdx =900–10Px +0.2I+5Py –4Pz

Where Qdx is the quantity demanded of Kelewele,

Px is the price of Kelewele per lb,

I is the consumer income in Ghana Cedis,

Py and Pz are the prices of two goods that are related to Kelewele.

 a) Based on the demand function above, is

d) What is the equation of the demand curve if consumer incomes are GHȼ 40, the price of good Y is GHȼ 20 and the price of good Z is GHȼ 27? 1 Mark

e) Graph the demand function for Kelewele from d)

Now suppose the weekly supply function for Kelewele at UPSA campus

is QSx = -260 + 10Px – 2Pi

Where QSx is the quantity supplied of Kelewele and Pi is the price of inputs used in preparing Kelewele.

f) What is the supply function if input prices are GHȼ 20?

g) Graph the supply curve from f)


1
Expert's answer
2021-07-29T13:05:03-0400

Let there be two goods, good 1 and good 2. They are called substitutes when the price of one good positively impacts the QD of the other good. They are called complements when the price of one good negatively impacts the QD of the other good.


(a) When the price of Z rises, the QD of kelewele falls by 4 units. Since the price of good Z negatively impacts the QD of kelewele, kelewele and good Z are perfect complements. 

(d) Substituting the values of I, Py, and Pz in the equation of the demand curve;

"Qdx=900-10Px+0.2I+5Py-4Pz\\Qdx=900-10Px+0.2(40)+5(20)-4(27)\\Qdx=900-10Px+8+100-108\\Qdx=900-10Px"


e)

When Px is zero, Qdx is 900. So, the x-axis intercept is (900, 0). When Qdx is zero, Px is 90. In the figure below, AB is the required demand curve





f)

[a] With price = GHc20

QSx = -260 + 10Px – 2Pi

Qsx = -260 +10Px -2(20)

Qsx = -300 +10Px


g)